OpenStudy (anonymous):

Derivative of sin inverse 2^x+1/1+4^x

OpenStudy (anonymous):

\[\sin^{-1} \left( 2^x+1\div4^x+1\right)\]

OpenStudy (anonymous):

please help :'(

OpenStudy (kainui):

Alright, well do you know the derivative of just y=sin^-1(x) ? After that, you're just applying the chain rule.

OpenStudy (anonymous):

can u please solve it... i cnt do it :(

OpenStudy (anonymous):

?

OpenStudy (anonymous):

ny1 ?

OpenStudy (kainui):

I mean if you can't do it, you're not gonna have a fun time on the test. I won't be taking your test for you.

OpenStudy (anonymous):

dude...plz...its kinda urgent :'(

zepdrix (zepdrix):

Kainui already asked this but you didn't answer for some reason. Do you know the derivative of \(\large sin^{-1} x\) ?

OpenStudy (anonymous):

yes... 1/under root1+x^2

OpenStudy (anonymous):

i tried it lyk 10 times... please help :/ i did apply the chain rule..still m not gettin the solution

zepdrix (zepdrix):

\[\large \frac{d}{dx}\sin^{-1}\color{orangered}{x} \qquad=\qquad \frac{1}{\sqrt{1-\color{orangered}{x}^2}}\] I think it's subtraction in the bottom, but ok that's a good start :) Let's apply that to our problem.

zepdrix (zepdrix):

\[\large \frac{d}{dx}\sin^{-1}\color{orangered}{\left(\frac{2^x+1}{4^x+1}\right)} \qquad=\qquad \frac{1}{\sqrt{1-\color{orangered}{\left(\dfrac{2^x+1}{4^x+1}\right)}^2}}\color{royalblue}{\left(\frac{2^x+1}{4^x+1}\right)'}\] Follow along so far? The blue term shows up due to the chain rule. The prime is to show us that we still need to take the derivative of that part.

OpenStudy (anonymous):

yes i got it.. now what?

zepdrix (zepdrix):

Now we have to apply the `Quotient Rule` to the blue part.

zepdrix (zepdrix):

Here is the setup for the quotient rule, \[\large \left(\frac{2^x+1}{4^x+1}\right)'=\frac{\color{royalblue}{(2^x+1)'}(4^x+1)-(2^x+1)\color{royalblue}{(4^x+1)'}}{(4^x+1)^2}\] The blue terms are the ones we have to take derivatives of. Understand how that was setup?

OpenStudy (anonymous):

yes..now ?

zepdrix (zepdrix):

Do you remember your exponential derivatives? They're a lil tricky. What's the derivative of \(\large 2^x+1\) ?

OpenStudy (anonymous):

2^x / log2

zepdrix (zepdrix):

mmm not division :) just multiplication, but yes good. \[\large (2^x)'=2^x(\ln2)\]

OpenStudy (anonymous):

now?

zepdrix (zepdrix):

\[\large \frac{\color{royalblue}{(2^x+1)'}(4^x+1)-(2^x+1)\color{royalblue}{(4^x+1)'}}{(4^x+1)^2}\] So taking the derivative of that first blue term gives us,\[\large \frac{\color{orangered}{(2^x(\ln2)+0)}(4^x+1)-(2^x+1)\color{royalblue}{(4^x+1)'}}{(4^x+1)^2}\] What do you get for the other blue part?

OpenStudy (anonymous):

4^x log 4

zepdrix (zepdrix):

\[\large \frac{\color{orangered}{2^x(\ln2)}(4^x+1)-(2^x+1)\color{orangered}{4^x(\ln4)}}{(4^x+1)^2}\] Ok good, let's put this together with our first part that we did. \[\frac{d}{dx}\sin^{-1}\color{orangered}{\left(\frac{2^x+1}{4^x+1}\right)} =\frac{1}{\sqrt{1-\color{orangered}{\left(\dfrac{2^x+1}{4^x+1}\right)}^2}}\color{royalblue}{\left(\frac{2^x+1}{4^x+1}\right)'}\] So we did the blue part, giving us,\[\frac{d}{dx}\sin^{-1}\color{orangered}{\left(\frac{2^x+1}{4^x+1}\right)} =\frac{1}{\sqrt{1-\color{orangered}{\left(\dfrac{2^x+1}{4^x+1}\right)}^2}}\color{royalblue}{\left(\large \frac{2^x(\ln2)(4^x+1)-(2^x+1)4^x(\ln4)}{(4^x+1)^2}\right)}\]

zepdrix (zepdrix):

ah it got cut off.. sec ill paste it again,

OpenStudy (anonymous):

ya sure

zepdrix (zepdrix):

\[\large \frac{1}{\sqrt{1-\color{orangered}{\left(\dfrac{2^x+1}{4^x+1}\right)}^2}}\color{royalblue}{\left(\large \frac{2^x(\ln2)(4^x+1)-(2^x+1)4^x(\ln4)}{(4^x+1)^2}\right)}\]

zepdrix (zepdrix):

I mean.. I guess you could simplify ... maybe.. But I don't see the point. That's your answer right there.

OpenStudy (anonymous):

thankyou :)

zepdrix (zepdrix):

np c: